OpenStudy (anonymous):

If a quarter of the weight of a roast is lost in roasting, what weight of roast should be bought in order to have 3kg of roasted meat?

OpenStudy (anonymous):

the answer is 4 Kgs

OpenStudy (anonymous):

Is it? thats what I think

OpenStudy (anonymous):

can you please explain how you worked it out?

OpenStudy (anonymous):

Ok think it this way: if 4kg is split into quarters, it would be 1kg each, if one quarter is taken away then it is taking away 1kg, 4-1=3kg

OpenStudy (anonymous):

Btw, what grade is this?

OpenStudy (anonymous):

That means I got it right before! THANK YOU and this is yr 9 work (:

OpenStudy (apoorvk):

okay so let the initial weight be 'x'. you lose 1/4th of the weight. so you lose x/4 so you are left with final minus initial x - x/4 now what is left 3 kg so, 3 = x- x/4 or, 3 = 3x/4 or, 12 = 3x or, x =4 so 4 kg is the initial weight of roast you need to take!